### v1

# Jialiang Fan Franklin College of Arts and Sciences

Abstract
Intracerebral hemorrhage (ICH) is a severe and sudden medical condition caused by the rupture of blood vessels in the brain, leading to permanent damage to brain tissue and often resulting in functional disabilities or death in patients. Diagnosis and analysis of ICH typically rely on brain CT imaging. Given the urgency of ICH conditions, early treatment is crucial, necessitating rapid analysis of CT images to formulate tailored treatment plans. However, the complexity of ICH CT images and the frequent scarcity of specialist radiologists pose significant challenges. Therefore, we built a dataset for ICH and normal classification and three types of ICH image classification based on the hemorrhage location, i.e., Deep, Subcortical, and Lobar. In addition, we propose a dual-task vision transformer (DTViT) for the automated classification and diagnosis of ICH images. This neural network utilizes the encoder from ViT, employing attention mechanisms for feature extraction from CT images. We incorporated two multilayer perception (MLP)-based decoders within the network to simultaneously identify the presence of ICH and classify three types of hemorrhage locations. Experimental results demonstrate that our proposed multi-classification network performs well on the built real-world test dataset. The code and dataset for this study will be made publicly available upon paper acceptance at: https://github.com/Jialiangfan/ICH-classification.
Keywords Intracerebral hemorrhage classification, vision transformer, dual-task classification.
1 Introduction
Intracerebral hemorrhage (ICH) is type of a severe condition characterized by the formation of a hematoma within the brain parenchyma. Representing 10-15% of all stroke cases, ICH is linked to significant morbidity and mortality rates[1]. Head computerized tomography (CT) is the standard method to diagnose ICH that can obtain accurate images of the head anatomical structure and detect abnormalities. Early ICH classification can provide appropriate early-stage treatment such as blood pressure management, hemostatic treatment, and surgical treatment. However, skilled doctors are often in very short supply in some regions, and they sometimes face a high volume of cases. Additionally, the location of intracerebral hemorrhage (ICH) is crucial for medical professionals when determining treatment strategies. Different types of ICH, such as lobar hemorrhage, subtentorial hemorrhage, and deep hemorrhage, present distinct challenges. Diagnosing the correct type of ICH can be particularly challenging even for experienced radiologists.
In recent years, deep-learning-based computer vision methods have been playing an important role in CT imaging diagnosing. For example, Sahu et al. [2] present an automated deep learning-based system for COVID-19 detection using chest CT images, achieving a 95.73% accuracy through advanced preprocessing and the proposed model, highlighting its efficiency over manual diagnostic methods. Huo et al. [3] present a general-purpose model for medical image classification that uses hierarchical multi-scale feature fusion to efficiently extract and integrate local and global features, significantly enhancing lesion detection in medical imagery. Huynh et al. [4] introduces a loss function named adaptive alended consistency loss for semi-supervised medical image learning that addresses class imbalance by blending class distributions, demonstrating superior performance over existing methods in experiments on imbalanced datasets. For the detection and classification of cerebral hemorrhage, numerous studies have utilized computer vision and deep learning techniques to aid medical professionals and enhance diagnostic efficiency[5, 6, 7, 8]. In literature [5], a deep learning model based on Resnet and EfficientDet that detects bleeding in CT scans is proposed, offering both classification and region-specific decision insights. Literature [6] demonstrates the efficacy of CNN-based deep learning models, particularly CNN-2 and ResNet-50, in classifying strokes from CT images, with future plans to optimize these models for improved diagnostic accuracy and efficiency. Li et al.[7] presents a rapid method using microwave signals to detect and classify cerebral hemorrhages, achieving high accuracy through a novel signal processing algorithm combined with neural networks, and demonstrating potential for quicker, life-saving diagnostics in emergency scenarios. However, these studies primarily focus on the task of classifying whether there is cerebral hemorrhage, but neglect further detection and diagnosis, i.e., the location and depth of the hemorrhage, which is more clinically significant.
Based on the above discussion, we propose a vision transformer (ViT)-based neural network model that achieves both ICH classification and hematoma location classification simultaneously.
2 Materials and methods
In this section, we introduce the dataset, the data preprocessing of ICH images, the construction of the proposed neural network model based on ViT, and the evaluation index. The structure of the proposed neural network is shown in Fig. 1.
2.1 Dataset
The dataset is sourced from the Department of Neurology at The First Hospital of Yulin. It includes 15,936 CT slices from 249 patients with intracerebral hemorrhage (ICH) collected between 2018 and 2021, and 6,445 CT slices from 199 healthy individuals from 2024. The healthy subjects have one set of CT images, while the ICH patients have two sets, captured within 24 hours and within 72 hours of symptom onset. All images were obtained using the LightSpeed VCT scanner from GE, USA, at Yulin First Hospital. All scans are saved as DCM files, featuring a resolution of 512×512 pixels, a slice thickness of 5 mm, and an inter-slice gap of either 5 mm or 1 mm.
In addition, each group of ICH images are manually classified into three different types according to hemorrhage location by physician: Deep ICH, Lobar ICH, and subtentorial ICH. To protect patient’s privacy, the dataset only includes patients’ gender and CT images. A detailed data description is listed in Table I.
Table 1: Data distribution of the dataset
Attribute Total Value Number Percentage
Sex $221$ Male $117$ $52.94\%$
Female $104$ $47.06\%$
ICH $12651$ Yes $8244$ $65.16\%$
No $4407$ $34.84\%$
Hemorrhage Location 8244 Deep $6093$ $73.91\%$
Lobar $1656$ $20.09\%$
Subtentorial $495$ $6\%$
Figure 1: CT image filtering. (a) The original CT image with a fixation device. (b) Processed image with fixation device removed.
2.2 Data preprocessing
The data preprocessing is composed of three stages: morphological treatment, manual filtering, and data augmentation.
2.2.1 Morphological treatment
Firstly, the pixel array is extracted from the DICOM file. Then, contrast enhancement and normalization are applied to the pixel array, which is finally converted to 8-bit integers. Additionally, when taking CT images, patients are equipped with fixation braces to keep their heads steady, which is also captured by the scanner and may add noise to the image data. Therefore, it is necessary to remove the artifact to purify the data. We binarized the image to highlight key features and applied erosion with a disk-shaped element to reduce noise. Then, Edge columns were zeroed out to eliminate potential artifacts, and flood filling was used to enhance specific areas. Finally, we superimposed the processed binary image onto one of the windowed images, thereby enhancing the visibility of regions of interest and improving the diagnostic utility of the scans. An instance of the original CT image and the processed result is displayed in Fig. 1.
2.2.2 Manual filtering
The head CT scan usually starts from the base of the brain (near the neck) and covers the entire brain up to the forehead. It means that only part of CT scans can capture the ICH of the patient and present a clear view for diagnoses, and other CT images may be similar between normal people and ICH patients. Therefore, each group of ICH CT images is manually selected by medical specialists to remove meaningless images to refine the dataset and reduce noise.
2.2.3 Data augumentation
As presented in Table 1, The number of images for three types of cerebral hemorrhage locations varies significantly. Therefore, we take data augumentation operation on subtentorial type CT images. Firstly, we augmented the dataset of three types of cerebral hemorrhage CT images proportionally to ensure that the distribution was approximately 1:1:1. Secondly, we also augmented the images without cerebral hemorrhage to achieve a roughly 1:1 ratio between the two categories.
As introduced in Section 2.1, we conduct experiments on the dataset we made from the real world. The augmented dataset is composed of a total of 30222 images. We divide the dataset as training, validating, and testing according to a ratio of 8:1:1. Furthermore, to enhance the diversity of the dataset, we applied image transformation on the training dataset, including center cropping to the size of 224x224, random rotation of 15 degrees, random sharpness adjustment with a factor of 2, and normalization with the mean value $[0.485,0.456,0.406]$ and standard deviation $[0.229,0.224,0.225]$.
Figure 2: CT image filtering. (a) The original CT image with a fixation device. (b) Processed image with fixation device removed.
2.3 Methods
This section introduces the methodology used in the paper. We discuss the ViT first, present the architecture of the developed neural network model, and introduce the evaluation index.
2.3.1 Vision Transformer
The Transformer, a network architecture based on the self-attention mechanism, has achieved remarkable success in the field of natural language processing (NLP) in recent years. Therefore, Researchers have extended the transformer model to the field of computer vision by segmenting images into patches and feeding them into the Transformer architecture. This approach computes attention weights between different pixel blocks, facilitating effective feature extraction from images. Experimental results demonstrate that Vision Transformers achieve superior and more promising performance compared to traditional CNN-based neural networks.
In ViT, a raw 2D image is initially transformed into a sequence of 1D patch embeddings. This process mirrors the word embedding technique used in natural language processing. In our research, the 2D CT image patch, denoted as $v$ is reshaped as a sequence of $N$ patches, where $P$ is the patch size set as $16\times 16$. These patches are then linearly projected into vectors of dimension.
In addition, to preserve the spatial context of the CT image patches, a trainable positional embedding is incorporated. This adjustment is essential due to the permutation invariance of self-attention mechanisms.
2.3.2 Model structure
Considering the high-density shadows, variable scales, and diverse locations characteristic of ICH CT images, we adopt the Vision Transformer (ViT) as the encoder of our neural network. Additionally, two decoders have been integrated into the network to achieve multi-task classfication. The decoders share the feature extraction results from the ViT encoder applied to the CT images. The extracted image features are then fed into two fully connected layers to classify the images. The structure of the NN model is shown in Figure 2.
The proposed neural network model is compose of an encoder for image feature extraction and two decoders that can simultaneously classify CT images, identifying whether there is a brain hemorrhage and the type of brain hemorrhage.
2.3.3 Evaluation indices
We have taken various experiments and adopt several models to test the performance. The correct classification of images is the criterion of the model performance, which is presented as
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN},$ (1)
where $TP$ represents true positive instances; $TN$ represents true negative instances; $FP$ represents false positive instances; $FN$ represents false negative instances. The precision metric represents the proportion of correctly predicted positive instances out of all the instances predicted as positive in a class. Precision is determined using the following equation:
$Precision=\frac{TP}{TP+FP}.$ (2)
The recall metric quantifies the proportion of positive instances that are correctly identified and the corresponding formula to calculate recall is presented as
$Recall=\frac{TP}{TP+TN}.$ (3)
The sensitivity metric calculates the proportion of positive instances that are correctly classified, which is defined as
$Sensitivity=\frac{TP}{TP+TN}$ (4)
In addition, the F1-score metric is the harmonic mean of precision and recall. It is computed as
$F1=\frac{2*(Precision*Recall)}{Precision+Recall}.$ (5)
We take these indices to evaluate our model’s performance.
Figure 3: The accuracy and loss curves in the training and validating process on datasets with and without augmentation. (a) The training and validating losses with augmentation. (b) The training and validating accuracy with augmentation. (c) The training and validating losses without augmentation. (d) The training and validating accuracy without augmentation.
3 Experiments
In this section, we conduct experiments using the built dataset with and without data augmentation to validate the efforts of the augmentation. Performance indexes introduced in Section 2.3.3 are utilized for evaluation. In addition, we also select several classical neural network models as encoders for feature extraction and conduct comparative experiments to assess their performance.
3.1 Environments and parameters
The experimental hardware setup is as follows: The CPU used is an AMD Ryzen 5975WX with 32 cores, and the GPU is an NVIDIA RTX A6000 equipped with 48 GB of memory. The operating system is Ubuntu 20.04. CUDA version 12.2 was utilized for computation. All experiments were conducted using Python version 3.8.18 and PyTorch version 2.2.1.
3.2 DTViT performances
Initially, we trained the model on both the original dataset and the augmented dataset, setting batch sizes of 8 and 32, respectively. We configured the training for 50 epochs, using Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.01. Figure 3 illustrates the evolution of loss and accuracy during the training and validation processes, both with and without data augmentation. As depicted in Figure 3(a), both training and validation losses decrease steadily throughout the training, eventually stabilizing at approximately 0.1. Correspondingly, both training and validation accuracies show steady increases, converging closely by the end of the training period. In contrast, results from the non-augmented dataset display inconsistency: while training loss decreases continuously, validation loss fluctuates and fails to converge, ending above 0.2 and indicating overfitting. Similarly, validation accuracy is unstable compared to training accuracy, with its final value significantly lower, suggesting poor model performance on unseen data.
Further, we apply the trained model on the testing dataset to get the performance, which is composed of 3023 CT images, including 1342 for normal images and 1681 ICH images with Deep 606, Lobar 479, and Subtentorial 596. The confusion matrix is shown in Fig. 4, where Fig. 4(a) shows the result of Classifier 1 for normal and ICH patients classification, and Fig. 4(b) shows the results of Classifier 2 for three types of ICH classification. It can be seen that the true negative cases take 1331 over 1332, and the true positive cases take 1691 over 1691 cases, achieving $99.23\%$ of the accuracy. Similarly, the classification of ICH types also achieved desirable results with a testing accuracy of $97.37\%$. In addition, the detailed data, including precision, recall, F1 score, and specificity, is displayed in Table 2.
Table 2: Performance comparison of DTViT with and without data augmentation on the testing data
Classifier DA Accuracy Precision Recall F1 Score Specificity
Classifier 1 Yes 0.992 0.996 0.987 0.991 0.997
Classifier 1 No 0.974 0.953 0.991 0.972 0.960
Classifier 2 Yes 0.974 0.962 0.969 0.966 0.995
Classifier 2 No 0.877 0.756 0.987 0.856 0.986
Figure 4: The confusion matrixes of two classifiers on the testing dataset. (a) The confusion matrix of classifier 1 for normal and ICH classification. (b) The confusion matrix of classifier 2 for three types of ICH classification.
3.3 Comparative experiments
We conducted comparative experiments using classical computer vision models, including VGG-16 [9] and ResNet18 [10], on our augmented dataset. The results, presented in Table 3, indicate that both models perform exceptionally well on the testing dataset. Specifically, ResNet18 achieves an accuracy of 0.999 on Classifier 1, while VGG16 reaches perfect accuracy, correctly classifying all test images on Classifier 1. The performance on Classifier 2, though slightly lower, remains impressive: ResNet18 and VGG16 achieve accuracies of 0.997 and 0.995, respectively. Overall, both ResNet18 and VGG16 slightly outperform the Vision Transformer (ViT) model.
Table 3: Performance comparison of two classifiers with and without data augmentation
Classifier Encoder Accuracy Precision Recall F1 Score Specificity
Classifier 1 ViT 0.989 0.997 0.979 0.988 0.998
Classifier 1 ResNet18 0.999 1 0.998 0.999 1
Classifier 1 VGG16 1 1 1 1 1
Classifier 2 ViT 0.971 0.97 0.967 0.968 0.988
Classifier 2 ResNet18 0.997 1 0.993 0.996 1
Classifier 2 VGG16 0.995 1 0.988 0.994 1
4 Discussion
In this paper, we first collect CT images from intracerebral haemorrhage (ICH) patients and normal people, which are sourced from real-world patient data from Yulin First Hospital, China. Furthermore, medical specialists put effort into data filtering and labelling of three types of ICH haemorrhage.
Based on the built dataset, we propose the dual-task vision transformer (DTViT) model based on the vision transformer for dual-task classification. The proposed model is composed of an encoder to extract information from CT images and two decoders for different classification tasks, i.e., classification of normal and ICH images and classification of three types of ICH based on the location of the hematoma. The proposed DTViT achieves 99.7% of the training accuracy and 99.5% of the testing accuracy on the augmented dataset. We also compared the proposed model with several state-of-the-art models, i.e., Resnet18, Resnet34, Vgg16, and Alexnet, and the results show that the proposed model achieves a desirable result.
There is some previous research focusing on ICH classification or detection[7, 5, 6, 8]. For example, literature[7] proposes an ICH classification and localization method using a neural network model, achieving the accuracy of 97.4%, while the input signals are microwave signals and the hardware requirements are relatively high. Also, a CT-image-based deep learning method is proposed in [5], which is based on the EfficientDet and achieves an accuracy of 92.7%. Similar work using computer vision methods for ICH detection and classification includes [11, 12], where literature [11] achieves ICH classification using a CNN-based architecture called EfficientNet; literature [12] uses ResNet-18 for ICH classification with the accuracy 95.93%.
However, to the best of our knowledge, there is neither such a dataset describing the location of hematoma of ICH CT images nor such classification model simultaneous determining whether a CT image is of a cerebral hemorrhage or normal, and classifies the three types of cerebral hemorrhage.
The primary aim of creating the Intracerebral Hemorrhage (ICH) CT image dataset and developing the classification model is to harness state-of-the-art computer vision technology to assist physicians in diagnosing and treating patients with cerebral hemorrhage. Cerebral hemorrhage is a severe, acute medical condition affecting approximately two million individuals annually, often associated with higher mortality and morbidity and limited treatment options. Early diagnosis and tailored treatment strategies based on the hemorrhage location are crucial for patient outcomes, as different hemorrhage sites require varied treatment approaches. Therefore, the classifier developed in this study, which categorizes hemorrhage based on its location, is clinically significant as it aids healthcare professionals in quickly and accurately determining treatment plans.
Limitations: This study faces several limitations. First, the dataset, derived from real clinical data, is challenging to obtain and inherently imbalanced, which may limit the model’s accuracy and generalizability. Future efforts will focus on expanding and balancing the dataset through continued data collection. Second, the current data augmentation technique involves mere replication of the dataset. Future improvements will explore the use of generative models, such as diffusion models, for data enhancement.
Future Directions: In subsequent work, we aim to develop a multimodal diagnostic dataset for cerebral hemorrhage, integrating clinical data such as blood pressure, lipid profiles, and bodily element levels to enhance diagnostic accuracy. Furthermore, we plan to create a cerebral hemorrhage classification and diagnosis model based on Artificial Intelligence-Generated Content (AIGC), which will improve diagnostic efficiency for physicians.
5 Conclusion
In this paper, we have constructed a dataset for Intracerebral Hemorrhage (ICH) based on real-world clinical data. The dataset comprises CT images that have been initially processed and manually classified by medical experts into categories of normal and ICH images. Moreover, the ICH images have been categorized into three types-Deep, Subcortical, and Lobar-according to the hemorrhage’s location.
Additionally, we have introduced a dual-task vision Transformer (DTViT) aimed at the classification of Intracerebral hemorrhage. This innovative neural network model incorporates an encoder, which utilizes the cutting-edge vision Transformer architecture, and two decoders that are designed to classify images as ICH or normal and to determine the hemorrhage type. Our experiments have demonstrated that the DTViT has achieved a remarkable accuracy rate of 99.7% on the test data. We have also compared the DTViT with traditional models such as ResNet, VGG, and AlexNet.
To our knowledge, this research is the first to have developed both a specialized dataset and a neural network model tailored for hemorrhage location classification in a clinical context. This contribution holds substantial potential for enhancing clinical diagnosis and treatment planning.
Acknowledgments
This was supported in part by the University of Georgia, and in part by the Department of Neurology, The First Hospital of Yulin.

### v2

@@
-Jialiang Fan Franklin College of Arts and Sciences
+Dual-Task Vision Transformer for Rapid and Accurate Intracerebral Hemorrhage Classification on CT Images
 Abstract
@@
 However, the complexity of ICH CT images and the frequent scarcity of specialist radiologists pose significant challenges.
-Therefore, we built a dataset for ICH and normal classification and three types of ICH image classification based on the hemorrhage location, i.e., Deep, Subcortical, and Lobar.
-In addition, we propose a dual-task vision transformer (DTViT) for the automated classification and diagnosis of ICH images.
-This neural network utilizes the encoder from ViT, employing attention mechanisms for feature extraction from CT images.
-We incorporated two multilayer perception (MLP)-based decoders within the network to simultaneously identify the presence of ICH and classify three types of hemorrhage locations.
-Experimental results demonstrate that our proposed multi-classification network performs well on the built real-world test dataset.
-The code and dataset for this study will be made publicly available upon paper acceptance at: https://github.com/Jialiangfan/ICH-classification.
-Keywords Intracerebral hemorrhage classification, vision transformer, dual-task classification.
+Therefore, we build a dataset from the real world for ICH and normal classification and three types of ICH image classification based on the hemorrhage location, i.e., Deep, Subcortical, and Lobar.
+In addition, we propose a neural network structure, i.e., dual-task vision transformer (DTViT), for the automated classification and diagnosis of ICH images.
+The DTViT utilizes the encoder from the Vision Transformer (ViT), employing attention mechanisms for feature extraction from CT images.
+Also, it incorporates two multilayer perception (MLP)-based decoders to simultaneously identify the presence of ICH and classify three types of hemorrhage locations.
+Experimental results demonstrate that DTViT performs well on the built real-world test dataset.
+The code and dataset for this work will be made publicly available upon paper acceptance at: https://github.com/Jialiangfan/ICH-classification.
+Keywords Intracerebral hemorrhage, image classification, vision transformer, transfer learning.
 1 Introduction
-Intracerebral hemorrhage (ICH) is type of a severe condition characterized by the formation of a hematoma within the brain parenchyma.
+Intracerebral hemorrhage (ICH) is a type of severe condition characterized by the formation of a hematoma within the brain parenchyma[1, 2].
 Representing 10-15% of all stroke cases, ICH is linked to significant morbidity and mortality rates[3].
 Head computerized tomography (CT) is the standard method to diagnose ICH that can obtain accurate images of the head anatomical structure and detect abnormalities.
-Early ICH classification can provide appropriate early-stage treatment such as blood pressure management, hemostatic treatment, and surgical treatment.
-However, skilled doctors are often in very short supply in some regions, and they sometimes face a high volume of cases.
-Additionally, the location of intracerebral hemorrhage (ICH) is crucial for medical professionals when determining treatment strategies.
-Different types of ICH, such as lobar hemorrhage, subtentorial hemorrhage, and deep hemorrhage, present distinct challenges.
-Diagnosing the correct type of ICH can be particularly challenging even for experienced radiologists.
-In recent years, deep-learning-based computer vision methods have been playing an important role in CT imaging diagnosing.
+However, analyses of head CT images for ICH classification and diagnosis usually require skilled radiologists, who are in very short supply in developing countries or regions.
+This may lead to delays in setting appropriate treatment plans and interventions, which are very crucial for accurate ICH patients.
+Furthermore, the diagnosis and analysis of ICH in CT images often require extensive experience and concentrated attention.
+However, when physicians are overworked or face a high volume of cases, misdiagnoses can inevitably occur.
+Therefore, the use of assistive medical technologies to aid physicians in diagnosing and analyzing ICH images can enhance the speed of diagnosis and treatment.
+This not only alleviates the shortage of medical resources but also improves diagnostic efficiency and accuracy, which are of significant importance to both physicians and patients.
+In recent years, computer vision methods based on deep learning have played a crucial role in CT imaging diagnostics.
+Neural network models trained on extensive datasets of real-world data have demonstrated high accuracy and efficiency on medical images.
+These models can be applied to tasks such as classification[4, 5], segmentation[6, 7], or enhancement of medical images[8, 9].
 For example, Sahu et al.
@@
 [12] introduces a loss function named adaptive alended consistency loss for semi-supervised medical image learning that addresses class imbalance by blending class distributions, demonstrating superior performance over existing methods in experiments on imbalanced datasets.
-For the detection and classification of cerebral hemorrhage, numerous studies have utilized computer vision and deep learning techniques to aid medical professionals and enhance diagnostic efficiency[5, 6, 7, 8].
-In literature [5], a deep learning model based on Resnet and EfficientDet that detects bleeding in CT scans is proposed, offering both classification and region-specific decision insights.
-Literature [6] demonstrates the efficacy of CNN-based deep learning models, particularly CNN-2 and ResNet-50, in classifying strokes from CT images, with future plans to optimize these models for improved diagnostic accuracy and efficiency.
-Li et al.[7] presents a rapid method using microwave signals to detect and classify cerebral hemorrhages, achieving high accuracy through a novel signal processing algorithm combined with neural networks, and demonstrating potential for quicker, life-saving diagnostics in emergency scenarios.
-However, these studies primarily focus on the task of classifying whether there is cerebral hemorrhage, but neglect further detection and diagnosis, i.e., the location and depth of the hemorrhage, which is more clinically significant.
-Based on the above discussion, we propose a vision transformer (ViT)-based neural network model that achieves both ICH classification and hematoma location classification simultaneously.
+There are also numerous studies that have utilized computer vision and deep learning techniques for ICH image detection and classification to aid medical professionals and enhance diagnostic efficiency[13, 14].
+In literature [13], a deep learning model based on ResNet and EfficientDet that detects bleeding in CT scans is proposed, offering both classification and region-specific decision insights and achieving an accuracy of 92.7%.
+Literature [14] demonstrates the efficacy of convolutional neural network (CNN)-based deep learning models, particularly CNN-2 and ResNet-50, in classifying strokes from CT images, with future plans to optimize these models for improved diagnostic accuracy and efficiency.
+Li et al.[15] propose propose a Unet-based neural network model to detect hemorrhage strokes of CT images, achieving an accuracy of 98.59%.
+However, these studies primarily focus on the task of classifying whether there is cerebral hemorrhage, neglecting further detection and diagnosis, i.e., the location and depth of the hemorrhage, which is more clinically significant.
+The Transformer, a network architecture based on the self-attention mechanism, has achieved remarkable success in the field of natural language processing (NLP) in recent years[16].
+Building on this success, researchers have extended the Transformer to the field of computer vision, introducing Vision Transformer (ViT)[17].
+The ViT model segments images into patches and feeds them into the Transformer architecture.
+Then, computes attention weights between different pixel blocks, facilitating effective feature extraction from images.
+Experimental results demonstrate that Vision Transformers achieve superior and more promising performance compared to traditional CNN-based neural networks.
+Transfer learning is widely used in deep learning to reduce training time and improve training efficiencies.
+A pre-trained model is firstly built on a large benchmark dataset such as ImageNet[18], and common features are captured and saved in the weights of models.
+When applying the model to a new dataset or different tasks, we can train the model using task-specific datasets on top of the pre-trained model.
+Transfer learning also reduces the need for large labeled datasets, making deep learning on small or imbalanced datasets possible, which is particularly suitable for medical images.
+Building on the aforementioned discussions, we first build an image dataset from real-world sources that includes head images of both healthy individuals and ICH patients, categorized into three types—Deep, Lobar, and Subtentorial—based on the location of the hemorrhage.
+Furthermore, a dual-task Vision Transformer (DTViT) is designed to simultaneously classify images of normal individuals and ICH patients, as well as categorize three types of ICH based on the hemorrhage location.
+We conduct experiments using the built dataset and the proposed model, and the results show that our proposed model achieves superior testing accuracy of $99.88\%$.
+The contributions of this paper are summarized as follows:
+  • We have constructed a real-world ICH image dataset comprising CT images of both normal individuals and patients classified into three types of ICH, addressing issues of insufficient brain hemorrhage image datasets and the lack of datasets for brain hemorrhage location, which is of great significance to both medical and computer vision research.
+  • Built upon the constructed dataset, We have developed a deep learning model, i.e., dual-task Vision Transformer (DTViT), which is capable of performing dual-classification tasks simultaneously: distinguishing CT images of normal and ICH patients and identifying the type of ICH according to the hemorrhage location.
+  • Our comparative experiments show that the DTViT model achieves 99.88% accuracy, outperforming existing CNN models and demonstrating the high quality of our dataset.
+The remainder of this paper is structured as follows: Section 2 describes the dataset, data preprocessing techniques, and the architecture of DTViT.
+Section 3 details the experimental setup, evaluation metrics, and the results obtained.
+Section 4 discusses the limitations of this study and directions for future research.
+Finally, Section 5 concludes the paper.
 2 Materials and methods
-In this section, we introduce the dataset, the data preprocessing of ICH images, the construction of the proposed neural network model based on ViT, and the evaluation index.
-The structure of the proposed neural network is shown in Fig.
-1.
+In this section, we introduce the dataset, the data preprocessing of ICH images, and the construction of DTViT.
 2.1 Dataset
@@
 To protect patient’s privacy, the dataset only includes patients’ gender and CT images.
-A detailed data description is listed in Table I.
-Table 1: Data distribution of the dataset
+A detailed data description is listed in Table 1.
+Table 1: Data distribution of the dataset.
 Attribute Total Value Number Percentage
@@
 Subtentorial $495$ $6\%$
-Figure 1: CT image filtering. (a) The original CT image with a fixation device. (b) Processed image with fixation device removed.
+Figure 1: Morphological treatment of CT images.
 2.2 Data preprocessing
@@
 2.2.1 Morphological treatment
-Firstly, the pixel array is extracted from the DICOM file.
-Then, contrast enhancement and normalization are applied to the pixel array, which is finally converted to 8-bit integers.
-Additionally, when taking CT images, patients are equipped with fixation braces to keep their heads steady, which is also captured by the scanner and may add noise to the image data.
+When taking CT images, patients are equipped with fixation braces to keep their heads steady, which is also captured by the scanner and may add noise to the image data.
 Therefore, it is necessary to remove the artifact to purify the data.
-We binarized the image to highlight key features and applied erosion with a disk-shaped element to reduce noise.
-Then, Edge columns were zeroed out to eliminate potential artifacts, and flood filling was used to enhance specific areas.
-Finally, we superimposed the processed binary image onto one of the windowed images, thereby enhancing the visibility of regions of interest and improving the diagnostic utility of the scans.
+We apply morphological processing to remove the fixation brace and normalize the images, converting them from DICOM to JPG format.
+The steps are as follows.
+Initially, the pixel array is extracted from the DICOM file and duplicated twice.
+The first duplicate is used to create a mask, while the second is for generating the final result.
+We start by binarizing the pixel array to emphasize key features and apply erosion using a disk-shaped element to minimize noise.
+Next, the edge columns are zeroed out to remove potential artifacts, and flood filling is performed to enhance specific areas, thus producing the mask image.
+Subsequently, this mask is applied to the second, normalized pixel array.
+The resulting image is then converted to 8-bit integers and saved as a JPG file.
 An instance of the original CT image and the processed result is displayed in Fig.
 1.
+It can be seen that the fixed brace has been successfully removed and the normalized image is much clearer than the original.
 2.2.2 Manual filtering
 The head CT scan usually starts from the base of the brain (near the neck) and covers the entire brain up to the forehead.
-It means that only part of CT scans can capture the ICH of the patient and present a clear view for diagnoses, and other CT images may be similar between normal people and ICH patients.
-Therefore, each group of ICH CT images is manually selected by medical specialists to remove meaningless images to refine the dataset and reduce noise.
+It means that only part of CT scans can capture the hemorrhage location and present a clear view for diagnoses, and other CT images do not include effective information and are useless for classification.
+Therefore, medical specialists manually diagnose and select each group of CT images to remove meaningless images to refine the dataset and reduce noise.
+This step is critical to constructing the dataset.
 2.2.3 Data augumentation
 As presented in Table 1, The number of images for three types of cerebral hemorrhage locations varies significantly.
-Therefore, we take data augumentation operation on subtentorial type CT images.
+Therefore, we take data augmentation operation on subtentorial type CT images.
 Firstly, we augmented the dataset of three types of cerebral hemorrhage CT images proportionally to ensure that the distribution was approximately 1:1:1.
 Secondly, we also augmented the images without cerebral hemorrhage to achieve a roughly 1:1 ratio between the two categories.
-As introduced in Section 2.1, we conduct experiments on the dataset we made from the real world.
-The augmented dataset is composed of a total of 30222 images.
-We divide the dataset as training, validating, and testing according to a ratio of 8:1:1.
+The augmented dataset consists of 30,222 images in total, including 13,221 normal images and 17,001 ICH images, which are further categorized into 6,093 Deep, 5,940 Lobar, and 4,968 Subtentorial images.
 Furthermore, to enhance the diversity of the dataset, we applied image transformation on the training dataset, including center cropping to the size of 224x224, random rotation of 15 degrees, random sharpness adjustment with a factor of 2, and normalization with the mean value $[0.485,0.456,0.406]$ and standard deviation $[0.229,0.224,0.225]$.
@@
 2.3 Methods
-This section introduces the methodology used in the paper. We discuss the ViT first, present the architecture of the developed neural network model, and introduce the evaluation index.
-2.3.1 Vision Transformer
-The Transformer, a network architecture based on the self-attention mechanism, has achieved remarkable success in the field of natural language processing (NLP) in recent years.
-Therefore, Researchers have extended the transformer model to the field of computer vision by segmenting images into patches and feeding them into the Transformer architecture.
-This approach computes attention weights between different pixel blocks, facilitating effective feature extraction from images.
-Experimental results demonstrate that Vision Transformers achieve superior and more promising performance compared to traditional CNN-based neural networks.
-In ViT, a raw 2D image is initially transformed into a sequence of 1D patch embeddings.
-This process mirrors the word embedding technique used in natural language processing.
-In our research, the 2D CT image patch, denoted as $v$ is reshaped as a sequence of $N$ patches, where $P$ is the patch size set as $16\times 16$.
-These patches are then linearly projected into vectors of dimension.
-In addition, to preserve the spatial context of the CT image patches, a trainable positional embedding is incorporated.
-This adjustment is essential due to the permutation invariance of self-attention mechanisms.
-2.3.2 Model structure
-Considering the high-density shadows, variable scales, and diverse locations characteristic of ICH CT images, we adopt the Vision Transformer (ViT) as the encoder of our neural network.
-Additionally, two decoders have been integrated into the network to achieve multi-task classfication.
-The decoders share the feature extraction results from the ViT encoder applied to the CT images.
-The extracted image features are then fed into two fully connected layers to classify the images.
-The structure of the NN model is shown in Figure 2.
-The proposed neural network model is compose of an encoder for image feature extraction and two decoders that can simultaneously classify CT images, identifying whether there is a brain hemorrhage and the type of brain hemorrhage.
-2.3.3 Evaluation indices
-We have taken various experiments and adopt several models to test the performance. The correct classification of images is the criterion of the model performance, which is presented as
+This section introduces the construction of DTViT.
+Firstly, we introduce the basics of the Vision Transformer and the architecture of DTViT.
+Then, the transfer learning and details of the pre-trained model are introduced.
+An overall diagram is shown in Fig.
+2.
+2.3.1 Structure of DTViT
+In ViT, a raw 2D image is initially transformed into a sequence of 1D patch embeddings, which mirrors the word embedding technique used in natural language processing.
+Similarly, the positional embeddings are applied to the constructed patch embeddings to retain information about the location of each patch in the original image.
+Then, the sequences are fed into the encoder of the Transformer, which consists of multi-head self-attention layers and MLP layers, with normalization and residual connections.
+This enables the model to capture intricate connections among different patch blocks.
+Finally, the output is processed by the decoder, i.e., the full-connection layer or MLP layer, to output the classification results.
+The architecture of the ViT encoder is depicted in Fig.
+2.
+Considering the high-density shadows, variable scales, and diverse locations characteristic of ICH CT images, we adopt the ViT’s encoder to the DTViT.
+Additionally, two MLP-based decoders have been integrated into the DTViT as decoders to achieve dual-task classification.
+As shown in Fig.
+2, Two decoders share the feature extraction results from the encoder, where the MLP classifier 1 achieves the classification of whether there is a hemorrhage in the image, and the MLP classifier 2 classifies the type of hemorrhage based on the location information extracted by the encoder.
+2.3.2 Transfer learning
+Transfer learning refers to training a model that has already been pre-trained on another large dataset, such as ImageNet[18].
+This approach can significantly save time and enhance training efficiency.
+It allows the model to leverage previously learned features and knowledge, which is especially useful when dealing with similar but new tasks.
+Therefore, we train the DTViT based on the pre-trained Vision Transformer encoder[19] and fine-tune the model’s parameters.
+Specifically, there are three types of pre-trained ViT models: ViT-Base, ViT-Large, and ViT-Huge.
+The ViT-Base model, with its 12 blocks, is suitable for datasets of moderate complexity.
+The ViT-Large model, which has 24 blocks and larger embedding dimensions, offers better performance but requires more computational resources.
+The ViT-Huge is the largest one of the three models, taking 32 blocks and being suitable for high-complex tasks.
+For our work with DTViT, we use the ViT-Large as the pre-trained model[20], which is trained on ImageNet-1K[18].
+The configuration of the pre-trained model is shown in Table 2.
+Table 2: Configuration and Parameters of ViT-Large Model.
+         Configuration          Value
+         Patch Size          16
+         Embedding Dimension          1024
+         Attention Heads          16
+         MLP Dimension          4096
+         Parameters          304326632
+3 Experiments
+In this section, we conduct experiments using the built dataset with and without data augmentation to validate the dataset and get the performance of the DTViT.
+The experiment environment and parameters used in the training process are first given.
+Then, we introduce several evaluation indices used in experiments.
+Further, we present the accuracies and losses of datasets.
+In addition, we apply the dataset to classical CNN models for comparison to depict the performance of our model.
+3.1 Environments and parameters
+The experimental hardware setup is as follows: The CPU used is an AMD Ryzen 5975WX with 32 cores, and the GPU is an NVIDIA RTX 4090 equipped with 24 GB of memory.
+The operating system is Ubuntu 20.04.
+CUDA version 12.3 was utilized for computation.
+All experiments were conducted using Python version 3.8.18 and PyTorch version 2.3.0.
+Additionally, we utilized the AdamW optimizer[21] with an initial learning rate of $2\times 10^{-5}$ and a weight decay of 0.01.
+The batch size is set as 8 for the training data without data augmentation, 32 for the training data with data augmentation, 32 for the validating data, and 4 for the testing data.
+The model was trained over 10 epochs.
+We employed the cross-entropy function as the loss function to optimize the model’s performance.
+3.2 Evaluation indices
+We have taken various experiments and adopted several models to get the performance of DTViT.
+Firstly, the correct classification of images, i.e., the accuracy, is the criterion of the model performance, which is presented as
 $Accuracy=\frac{TP+TN}{TP+TN+FP+FN},$ (1)
@@
 The sensitivity metric calculates the proportion of positive instances that are correctly classified, which is defined as
-$Sensitivity=\frac{TP}{TP+TN}$ (4)
+$Sensitivity=\frac{TP}{TP+TN}.$ (4)
 In addition, the F1-score metric is the harmonic mean of precision and recall. It is computed as
 $F1=\frac{2*(Precision*Recall)}{Precision+Recall}.$ (5)
-We take these indices to evaluate our model’s performance.
-Figure 3: The accuracy and loss curves in the training and validating process on datasets with and without augmentation.
-(a) The training and validating losses with augmentation.
-(b) The training and validating accuracy with augmentation.
-(c) The training and validating losses without augmentation.
-(d) The training and validating accuracy without augmentation.
-3 Experiments
-In this section, we conduct experiments using the built dataset with and without data augmentation to validate the efforts of the augmentation.
-Performance indexes introduced in Section 2.3.3 are utilized for evaluation.
-In addition, we also select several classical neural network models as encoders for feature extraction and conduct comparative experiments to assess their performance.
-3.1 Environments and parameters
-The experimental hardware setup is as follows: The CPU used is an AMD Ryzen 5975WX with 32 cores, and the GPU is an NVIDIA RTX A6000 equipped with 48 GB of memory.
-The operating system is Ubuntu 20.04.
-CUDA version 12.2 was utilized for computation.
-All experiments were conducted using Python version 3.8.18 and PyTorch version 2.2.1.
+In medical fields, specificity is a good evaluation index to guarantee that individuals who do not have the disease are not wrongly diagnosed, thereby avoiding unnecessary treatment and anxiety.
+The calculation of specificity is presented as
+$Specificity=\frac{TN+FP}{TN}.$ (6)
+We take these indices to evaluate the performances of DTViT on the built dataset with and without data augmentation.
 3.3 DTViT performances
-Initially, we trained the model on both the original dataset and the augmented dataset, setting batch sizes of 8 and 32, respectively.
-We configured the training for 50 epochs, using Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.01.
-Figure 3 illustrates the evolution of loss and accuracy during the training and validation processes, both with and without data augmentation.
-As depicted in Figure 3(a), both training and validation losses decrease steadily throughout the training, eventually stabilizing at approximately 0.1.
-Correspondingly, both training and validation accuracies show steady increases, converging closely by the end of the training period.
-In contrast, results from the non-augmented dataset display inconsistency: while training loss decreases continuously, validation loss fluctuates and fails to converge, ending above 0.2 and indicating overfitting.
-Similarly, validation accuracy is unstable compared to training accuracy, with its final value significantly lower, suggesting poor model performance on unseen data.
-Further, we apply the trained model on the testing dataset to get the performance, which is composed of 3023 CT images, including 1342 for normal images and 1681 ICH images with Deep 606, Lobar 479, and Subtentorial 596.
-The confusion matrix is shown in Fig.
+Figure 3 illustrates the evolution of losses and accuracies during training and validating processes, both with and without data augmentation.
+As depicted in Fig.
+3(a), training and validating losses decrease steadily throughout the process, eventually stabilizing at approximately 0.01.
+Correspondingly, both training and validating accuracies show steady increases, converging approximately to 0.997 by the end of the training period.
+By contrast, results from the non-augmented dataset display inconsistency: while training loss decreases continuously to a desirable value, the validating loss fails to converge to the same level, ending around 0.08 and indicating overfitting.
+Similarly, the validating accuracy does not reach the level of the training accuracy, suggesting poor performance on unseen data.
+Further, we evaluate the trained model on the test dataset to get the DTViT’s performance on the unknown data, which is composed of 1266 CT images, including 459 for normal images and 807 ICH images with Deep 595, Lobar 163, and Subtentorial 49.
+Confusion matrixes are shown in Fig.
 4, where Fig.
-4(a) shows the result of Classifier 1 for normal and ICH patients classification, and Fig.
-4(b) shows the results of Classifier 2 for three types of ICH classification.
-It can be seen that the true negative cases take 1331 over 1332, and the true positive cases take 1691 over 1691 cases, achieving $99.23\%$ of the accuracy.
-Similarly, the classification of ICH types also achieved desirable results with a testing accuracy of $97.37\%$.
-In addition, the detailed data, including precision, recall, F1 score, and specificity, is displayed in Table 2.
-Table 2: Performance comparison of DTViT with and without data augmentation on the testing data
+4(a) shows the result of Task 1 for normal and ICH patients classification, and Fig.
+4(b) shows the results of Task 2 for three types of ICH classification.
+It can be seen that almost all positive and negative cases on the test dataset are classified correctly, except 1 normal image is misclassified as an ICH image.
+Similarly, the classification of ICH types also achieved excellent results with a testing accuracy of $0.996$.
+In detail, evaluation indices of precision, recall, F1 score, and specificity are displayed in Table 3.
+Figure 3: The accuracy and loss curves in training and validating processes on datasets with and without augmentation.
+(a) Training and validating losses with augmentation.
+(b) Training and validating accuracies with augmentation.
+(c) Training and validating losses without augmentation.
+(d) Training and validating accuracies without augmentation.
+Table 3: Performance comparison of DTViT with and without data augmentation on the testing dataset.
 Classifier DA Accuracy Precision Recall F1 Score Specificity
-Classifier 1 Yes 0.992 0.996 0.987 0.991 0.997
-Classifier 1 No 0.974 0.953 0.991 0.972 0.960
-Classifier 2 Yes 0.974 0.962 0.969 0.966 0.995
-Classifier 2 No 0.877 0.756 0.987 0.856 0.986
-Figure 4: The confusion matrixes of two classifiers on the testing dataset.
-(a) The confusion matrix of classifier 1 for normal and ICH classification.
-(b) The confusion matrix of classifier 2 for three types of ICH classification.
+Classifier 1 Yes 1 1 1 1 1
+Classifier 1 No 0.996 0.991 1 0.995 0.995
+Classifier 2 Yes 0.996 0.998 0.997 0.997 1
+Classifier 2 No 0.992 0.994 0.996 0.995 0.995
+Figure 4: Confusion matrixes of DTViT for two classification tasks on the testing dataset.
+(a) Confusion matrix of task 1 for normal and ICH classification.
+(b) Confusion matrix of Task 2 for three types of ICH classification.
 3.4 Comparative experiments
-We conducted comparative experiments using classical computer vision models, including VGG-16 [9] and ResNet18 [10], on our augmented dataset.
-The results, presented in Table 3, indicate that both models perform exceptionally well on the testing dataset.
-Specifically, ResNet18 achieves an accuracy of 0.999 on Classifier 1, while VGG16 reaches perfect accuracy, correctly classifying all test images on Classifier 1.
-The performance on Classifier 2, though slightly lower, remains impressive: ResNet18 and VGG16 achieve accuracies of 0.997 and 0.995, respectively.
-Overall, both ResNet18 and VGG16 slightly outperform the Vision Transformer (ViT) model.
-Table 3: Performance comparison of two classifiers with and without data augmentation
-Classifier Encoder Accuracy Precision Recall F1 Score Specificity
-Classifier 1 ViT 0.989 0.997 0.979 0.988 0.998
-Classifier 1 ResNet18 0.999 1 0.998 0.999 1
-Classifier 1 VGG16 1 1 1 1 1
-Classifier 2 ViT 0.971 0.97 0.967 0.968 0.988
-Classifier 2 ResNet18 0.997 1 0.993 0.996 1
-Classifier 2 VGG16 0.995 1 0.988 0.994 1
+To better demonstrate the performance of the DTViT, we conducted comparative experiments using classical CNN models, including ResNet18[22], AlexNet[23], SqueezeNet[24], and DenseNet[25] on our augmented dataset.
+The comparative results are shown in Table 4.
+Specifically, the proposed model, i.e.
+DTViT, achieves the lowest loss of $0.0102$ and the highest accuracy of $0.9988$ over these models.
+In addition, the DTViT also ranks first on other evaluation metrics, except its precision is slightly lower than DenseNet’s.
+Table 4: Overall performance comparisons with existing models.
+Model Loss Accuracy Precision Recall F1 Score Specificity
+Proposed 0.0102 0.9988 0.9991 0.9983 0.9987 1
+ResNet18[22] 0.0170 0.9984 0.9989 0.9983 0.9986 0.9993
+AlexNet[23] 0.0142 0.9976 0.9989 0.9983 0.9986 0.9994
+SqueezeNet[24] 0.1244 0.9775 0.990 0.9889 0.9895 0.9993
+DenseNet[25] 0.0119 0.9976 1 0.9964 0.9981 1
 4 Discussion
@@
 The proposed model is composed of an encoder to extract information from CT images and two decoders for different classification tasks, i.e., classification of normal and ICH images and classification of three types of ICH based on the location of the hematoma.
-The proposed DTViT achieves 99.7% of the training accuracy and 99.5% of the testing accuracy on the augmented dataset.
-We also compared the proposed model with several state-of-the-art models, i.e., Resnet18, Resnet34, Vgg16, and Alexnet, and the results show that the proposed model achieves a desirable result.
+The proposed DTViT achieves 99.7% of the training accuracy and 99.88% of the testing accuracy on the augmented dataset.
+We also compared the proposed model with several state-of-the-art models, i.e., Resnet18, SqueezeNet, and Alexnet, and the results show that our model achieves the best performance over these models.
 There is some previous research focusing on ICH classification or detection[26, 13, 14, 27].
@@
 The dataset comprises CT images that have been initially processed and manually classified by medical experts into categories of normal and ICH images.
-Moreover, the ICH images have been categorized into three types-Deep, Subcortical, and Lobar-according to the hemorrhage’s location.
+Moreover, the ICH images have been categorized into three types, Subcortical, and Lobar, according to the hemorrhage’s location.
 Additionally, we have introduced a dual-task vision Transformer (DTViT) aimed at the classification of Intracerebral hemorrhage.
 This innovative neural network model incorporates an encoder, which utilizes the cutting-edge vision Transformer architecture, and two decoders that are designed to classify images as ICH or normal and to determine the hemorrhage type.
-Our experiments have demonstrated that the DTViT has achieved a remarkable accuracy rate of 99.7% on the test data.
-We have also compared the DTViT with traditional models such as ResNet, VGG, and AlexNet.
+Our experiments have demonstrated that the DTViT has achieved a remarkable accuracy rate of $99.88\%$ on the test data.
+We have also compared the DTViT with traditional models such as Resnet18, SqueezeNet, and Alexnet.
 To our knowledge, this research is the first to have developed both a specialized dataset and a neural network model tailored for hemorrhage location classification in a clinical context.